1. Water from Thonon, France has the following composition Anions mg/L HCO3 Soa2 NOa CI- Cations mg/L 332 14 14 8.2 103.2 16.1 1.4 5.1 Ca2+ 0A 2+ Mg a) What is the ionic strength of this water b) What are the activity coefficients for HCO3 and cO32 in this water? c) Assuming an equilibrium constant for the dissociation of bicarbonate of 4.68E-11 and a pH of 7.3, what is the equilibrium concentration of CO3 in this water?

What is the ionoc strength of the water?

What are the activity coefficients?

What is the equilibrium concentration of CO3-2
Answer #1

1. For the given data

concentration of,

[HCO3-] = 0.332 g/61.017 g/mol = 5.44 M

[SO4^2-] = 0.014 g/96 g/mol = 0.15 M

[NO3-] = 0.014 g/62 g/mol = 0.23 M

[Cl-] = 0.0082 g/35.4 g/mol = 0.23 M

a) Ionic strength = 1/2Cizi^2

                           = 1/2(5.44 + 0.15 x 4 + 0.23 + 0.23)

                           = 3.25

b) acitivty coefficient for,

log[Y(HCO3-)] = -0.51 x 1 x sq.rt.(3.25)/(1+3.3 x 0.45 x sq.rt.(3.25)

[Y(HCO3-)] = 0.56

[CO3^2-] = sq.rt.(4.8 x 10^-11 x 5.44) = 1.6 x 10^-5 M

New ionic strength = 0.53

log[Y(CO3^2-)] = -0.51 x 1 x sq.rt.(0.53)/(1+3.3 x 0.45 x sq.rt.(0.53)

Y(HCO3-) = 0.66

c) pH = -loh[H+] = 7.3

[H+] = 5.01 x 10^-8 M

4.68 x 10^-11 = 5.01 x 10^-8 [CO3^2-]/5.44

[CO3^2-] at equilibrium = 5.08 x 10^-3 M